Diana purchased 6 pounds of strawberries and 4 pounds of apples for $18.90. Then she realizes that was not enough and purchased 3 more pounds of each fruit for $10.74. What was the cost per pound for each type of fruit?
Answer by mananth(16946)

strawberries x $/lb , Price = $9.00
Apples y $/lb , Price = $55.00
Total 345
6 x + 4 y = 18.9 .............1
3 x + 3 y = 10.74 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 4
-18 x -12 y = -56.7
12 x + 12 y = 42.96
Add the two equations
-6 x = -13.74
/ -6
x = 2.29
plug value of x in (1)
6 x + 4 y = 18.9
13.74 + 4 y = 18.9
4 y = 18.9 -13.74
4 y = 5.16
y = 1.29
strawberries $/lb 2.29
Apples $/lb 1.29
m.ananth@hotmail.ca